# Recent Advances in Named Entity Recognition: A Comprehensive Survey and Comparative Study

### v2

@@ 8.1 Evaluation strategies: exact or relaxed evaluation
The evaluation of NER systems is based on comparing predictions with a gold standard, and typically employs one of two strategies: exact evaluation or relaxed evaluation.
  • Exact Evaluation: In this approach, both the boundaries and the type of the named entity must match the gold standard precisely. This stringent method requires a perfect alignment between the predicted entity and the reference and is commonly used in the CoNLL-2003 evaluation (Sang and De Meulder, 2003).
  • Relaxed Evaluation: This approach allows partial credit, giving points when either the type or the boundaries are correct, even if both are not. Relaxed evaluation is often used in the MUC (Grishman and Sundheim, 1996b) and ACE (Doddington et al., 2004) standards.
@@ 10.1 Datasets
Our experiments were conducted using ten datasets featuring texts from multiple domains, as detailed in Table 4. The datasets vary in size and number of classes, offering a broad evaluation across different scenarios.
@@
For GPT-4, custom prompts were created for each dataset. These prompts were tailored to outline the primary categories of named entities relevant to each dataset and included several examples. Figure 8 presents examples of prompts designed for the BC5CDR and MIT Restaurant datasets, accompanied by sample JSON-formatted responses.
Figure 8: Two examples of a prompt for detecting named entities: one for the BC5CDR dataset on the left and one for MITRestaurant on the right
@@ 10.2 Models
Our selection of frameworks was guided by three primary criteria: open-source availability, free access, and the ability to train models on custom datasets. Based on these factors, we included Flair, Stanford CoreNLP, spaCy, GliNER, OpenAI, and Hugging Face in our analysis. NLTK was excluded due to its reliance on the Stanford CoreNLP tagger, which uses a CRF model, thus introducing redundancy. For frameworks that support multiple algorithms, we selected several representative models.
For spaCy, we evaluated three models, each illustrating different architecture and capability levels: a small CNN (‘‘en_core_web_sm’’), a large CNN (‘‘en_core_web_lg’’), and a Transformer model based on basic RoBERTa (Liu et al., 2019a) (‘‘en_core_web_trf’’). In the case of Hugging Face, we chose six models: the basic BERT architecture in both lowercase and uppercase configurations (‘‘BERT-base-cased’’ and ‘‘BERT-base-uncased’’), its distilled variant (‘‘DistilBERT-base-cased“), a large RoBERTa model (‘‘FacebookAI/xlm-roberta-large’’), and both small and large versions of DeBERTa (‘‘Microsoft/DeBERTa-v3-base’’ and ‘‘Microsoft/DeBERTa-v3-large’’). For OpenAI, we used the GPT-4o model, while for GliNER, we selected the large architecture based on DeBERTa (“urchade/gliner_large”), both in its pre-trained and fine-tuned versions.
@@
For the Hugging Face models, we used the TNER library to fine-tune each model, with entity decoding managed by CRF (see Section 5.3.3). Hyperparameter tuning involved a grid search with learning rates of $10^{-4}$ and $10^{-5}$, both with and without weight decay set to 0.01. Training was conducted over 10 epochs, with an interim evaluation at 5 epochs, a batch size of 16, and gradient accumulation steps set to 2. Additionally, configurations included both CRF and non-CRF models, with a warmup step ratio of 0.1.
@@
Lastly, for GliNER, we fine-tuned the large model (“urchade/gliner_large”) on each dataset using customized hyperparameters. The main model was trained with a learning rate of $5\times 10^{-6}$ and weight decay set to 0.01. We employed a linear learning rate scheduler with a warmup ratio of 0.1. Training was conducted over 10 epochs with a batch size of 16, and focal loss parameters were set to alpha = 0.75 and gamma = 2. Step-based evaluations were performed every 300 steps.
@@ 10.3 Evaluation and statistical analysis
10.4 Results and discussion
@@
Table 5 provides a comparison of the selected NER frameworks based on their macro-averaged F1 scores across various datasets. The best and second-best performances are highlighted in bold and underlined, respectively. Our experiments reveal key performance trends influenced by dataset size, domain specificity, and the comparative effectiveness of pre-trained versus fine-tuned NER models. In the following, we delve into the details of these results and their implications for model performance across various datasets.
Starting with large general-domain datasets such as CoNLL-2003, OntoNotes, Few-NERD, and MultiCoNER, Transformer-based models such as GliNER-L (fine-tuned), RoBERTa, and DeBERTa demonstrate outstanding performance. This can likely be attributed to their high parameter counts, which enable them to capture complex data relationships when trained on extensive datasets. A key insight from this analysis is the notable impact of GliNER-L’s comprehensive pre-training and fine-tuning processes. Compared to DeBERTa from Microsoft, GliNER-L exhibits improved performance, especially on domain-specific datasets. This improvement is likely due to GliNER-L’s structured approach, which involves sequential stages: pre-training as a language model, followed by NER training, and fine-tuning on domain-specific datasets. These stages allow GliNER-L to capture nuanced entity relationships and effectively adapt to NER tasks.
Shifting the focus to domain-specific datasets, particularly in the biomedical field (BioNLP2004, NCBI Disease, and BC5CDR), a different trend emerges. Traditional models such as CRF and LSTM-CRF remain competitive, often closely matching or even surpassing Transformer-based models. This may be due to the specialized terminology in biomedical texts, which is not well represented in the general datasets typically used for Transformer pre-training. However, fine-tuning on domain-specific data, as demonstrated with GliNER-L, significantly enhances performance by refining embeddings to better capture the contextual nuances of specialized terms. This trend is illustrated in Table 6, which compares the performance of CRF, LSTM-CRF, GliNER-L (fine-tuned) and Microsoft’s DeBERTa on key types of entities in biomedical datasets.
@@
In the BioNLP2004 dataset, which covers a broader range of entity types (“cell_line”, “cell_type”, “DNA”, “Protein”, and “RNA”), CRF consistently achieves high scores, especially in precision across entities like “cell_type” (94.54), “DNA” (89.32), “Protein” (93.96), and “RNA” (93.0). For "Protein" and "RNA" entities, GliNER-L shows competitive recall values (87.13 and 81.36, respectively), indicating its strength in capturing complex biomedical terms despite not leading in all F1 scores. LSTM-CRF and DeBERTa display relatively balanced scores across categories, though GliNER-L has a slight edge in recall for multi-word biomedical terms.
@@
Turning to smaller datasets such as FIN and WNUT2017, we observe that LSTM-CRF models tend to outperform Transformer-based models. This difference may stem from the tendency of overparameterized models, such as Transformers, to overfit when trained on limited data. For instance, on the FIN dataset, LSTM-CRF achieves the highest F1 score (69.22), significantly outperforming Transformer models, which require careful hyperparameter tuning and well-structured validation sets to reduce overfitting. Notably, GliNER-L fine-tuned remains competitive even on these smaller datasets, consistently ranking among the top three models. This suggests that GliNER-L’s structured pre-training and fine-tuning processes enhance its robustness and adaptability across datasets of varying sizes.
@@
  • Precision/recall balance: Models positioned near the central line, such as DeBERTa variants and LSTM-CRF, exhibit a well-balanced trade-off between precision and recall. This balance makes them particularly suitable for tasks where an equilibrium between these metrics is crucial, ensuring both high accuracy and comprehensive coverage in entity detection.
  • Precision-focused models: Models positioned above the reference line, such as CRF and CNNs (en_core_web_sm and en_core_web_lg), demonstrate higher precision than recall, making them ideal for applications where minimizing false positives is a priority.
  • Recall-focused model: GliNER-L pre-trained model favors recall, demonstrating a considerably higher recall than precision, though it still falls significantly behind the other models in overall performance.
@@
Furthermore, Figure 10 presents a PCA plot (left) and hierarchical clustering (right), which illustrate performance similarities across frameworks. In the PCA plot, each point represents a model, with colors indicating clusters based on performance similarity. The first two dimensions capture most of the variance, showing that Transformer-based models, such as DeBERTa (base and large) and GliNER-L fine-tuned, cluster closely, suggesting consistent performance across datasets. In contrast, isolated points such as GPT-4o and GliNER-L pre-trained exhibit lower performance, particularly without fine-tuning. Hierarchical clustering (right) reinforces these insights. Traditional models, such as CRF and CNNs (both en_core_web_sm and en_core_web_lg), form a distinct cluster, highlighting their competitiveness but limited flexibility compared to Transformers. Fine-tuned Transformers, especially GliNER-L, stand out, underscoring the advantages of domain-specific training.
@@
Statistical analysis supports these findings. A Friedman test applied to all entity types in our datasets reveals statistically significant differences between model medians (p-value = 0.000), confirming the significance at an alpha level of 5%. The results of Nemenyi’s post hoc test (Figure 11) show that certain groups of models have similar performance levels, as indicated by the connecting horizontal lines. For instance, models such as GliNER-L fine-tuned and DeBERTa (large) achieve the best rankings, indicating superior performance, and are not significantly different from each other. This suggests that these models perform similarly on the evaluated datasets. Notably, among the Transformer-based models, the DeBERTa variants appear to be the most effective for NER.
Additionally, another cluster of models, including GliNER-L pre-trained, CNNs (en_core_web_sm and en_core_web_lg), and GPT-4o, also exhibit statistically similar performance, although with higher mean ranks, indicating relatively lower performance compared to the top-performing models. Toward the right of the figure, models such as CRF, DistilBERT, and BERT exhibit the highest ranks, indicating inferior performance in comparison to the other models.
@@
Our experiments have shown that while LLMs perform exceptionally well in many NLP tasks, their general nature often leads to a lack of specialized effectiveness for NER, which can impede precise entity recognition, particularly in noisy or dynamic contexts. Evaluations highlight a performance gap for NER tasks (Han et al., 2023), likely due to limited task-specific learning and explicit understanding, which may lead to a “lack of specialty” in NER. Moreover, the large, sometimes undisclosed, number of parameters in LLMs poses challenges for fine-tuning to meet the demands of NER. Consequently, directly addressing NER tasks using only LLMs remains difficult. However, combining LLMs with simpler, fine-tuned NER models offers a promising solution. Approaches such as LinkNER (Zhang et al., 2024b) leverage a hybrid model in which the NER component handles common entities, while the LLM, guided by uncertainty estimation, addresses complex or ambiguous cases, enhancing robustness and flexibility in open domain contexts. Similarly, the Super In-Context Learning (SuperICL) approach (Xu et al., 2023) integrates smaller, specialized plug-in models to provide task-specific predictions and confidence scores, refining LLM outputs for more accurate final predictions. Together, LinkNER and SuperICL demonstrate that combining the broad contextual knowledge of LLMs with the focused precision of smaller models offers a promising path to advance NER, particularly in challenging open-domain settings with unseen entities and noisy data.
@@ 11 Conclusion and perspectives
This article presents a comprehensive survey on recent advances in NER within a classification framework. We focus on recent methods, including LLMs, graph-based approaches, reinforcement learning techniques, and strategies for training models on small datasets. To assess these approaches, we evaluated popular frameworks across datasets with varying characteristics.
@@
Looking ahead, the potential of LLMs to enhance NER should not be underestimated. Despite current limitations in handling specialized NER tasks, the rapid advancement of language models presents an exciting opportunity to integrate these technologies into more refined NER systems. Promising directions for advancing NER include hybrid approaches that combine LLMs with specialized, fine-tuned NER models (such as LinkNER or SuperICL). By blending the broad contextual understanding of LLMs with the precision of specialized NER models, these strategies hold promise for tackling NER in dynamic, open-domain settings that involve unseen entities and noisy data.
Future research should further investigate these hybrid approaches, along with fine-tuning techniques and preprocessing methods designed to enhance LLM adaptability for NER. Such strategies could improve the applicability and accuracy of NER systems across diverse datasets and challenging contexts.

### v3

@@ 8.1 Evaluation strategies: exact or relaxed evaluation
The evaluation of NER systems is based on comparing predictions with a gold standard and typically employs one of two strategies: exact evaluation or relaxed evaluation.
  • Exact Evaluation: In this approach, both the boundaries and the type of the named entity must match the gold standard accurately. This stringent method requires a perfect alignment between the predicted entity and the reference and is commonly used in the CoNLL-2003 evaluation (Tjong Kim Sang and De Meulder, 2003).
  • Relaxed Evaluation: This approach allows partial credit, giving points when either the type or the boundaries are correct, even if both are not. Relaxed evaluation is often used in the MUC (Grishman and Sundheim, 1996) and ACE (Doddington et al., 2004) standards.
@@ 10.1 Datasets
Our experiments were carried out on ten datasets from various domains, as shown in Table 4. These datasets differ in size and class count, allowing for various evaluation scenarios.
@@
For GPT-4, custom prompts were designed for each dataset to highlight the key categories of named entities specific to that dataset. These prompts also included multiple examples to provide context. The box below showcases sample prompts for the BIONLP2004 dataset.
Example of a prompt used for the dataset BIONLP2004 Identify the named entities in the following sentence, categorizing them according to these definitions: "DNA": Entities that refer to specific DNA sequences or are related to DNA in a biological context. Examples: BRCA1 gene, CD14 5’-Upstream Sequence, or nonoptimal binding sites. "Protein": Entities representing specific proteins or related to proteins in a biological sense. Examples: c-myb, hemoglobin, or E-box-binding repressor. "Cell_type": Entities referring to specific types of cells in a biological context. Examples: neuron, T-cell, or hepatocyte. "Cell_line": Entities representing specific cell lines used in biological research. Examples: Monocytic U937 cells, Jurkat cell line, or MCF-7. "RNA": Entities that are specific RNA sequences or related to RNA in a biological context. Examples: mRNA, siRNA, or c-jun mRNA. Format the output in JSON with keys corresponding to each entity type. List the identified named entities under each key, grouping multiple entities of the same type into a single list. Ensure a clear separation and precise identification of each entity to avoid ambiguity. Return only the JSON response with no additional explanations.
@@ 10.2 Models
Our selection of frameworks was guided by three primary criteria: open-source availability, free access, and the ability to train models on custom datasets. Based on these factors, we included Flair Akbik et al. (2019), Stanford CoreNLP Manning et al. (2014), spaCy Honnibal and Montani (2017), GliNER Zaratiana et al. (2023), OpenAI OpenAI , and Hugging Face Wolf et al. (2020) in our analysis. For frameworks that support multiple algorithms, we selected several representative models.
For spaCy, we evaluated three models, each illustrating different architectures and capability levels: a small CNN (“en_core_web_sm”), a large CNN (“en_core_web_lg”), and a Transformer model based on basic RoBERTa Liu et al. (2019a) (“en_core_web_trf”).
In the case of Hugging Face, we chose six models: the basic BERT Devlin et al. (2018) architecture in both lowercase and uppercase configurations (“BERT-base-cased” and “BERT-base-uncased”), its distilled variant Sanh et al. (2019) (“DistilBERT-base-cased”), a large RoBERTa-based model (“FacebookAI/xlm-roberta-large”), and both small and large versions of DeBERTa He et al. (2020a) (“Microsoft/DeBERTa-v3-base” and “Microsoft/DeBERTa-v3-large”).
For OpenAI, we used the GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ model, while for GliNER, we selected the large architecture based on DeBERTa (“urchade/gliner_large”), both in its pre-trained and fine-tuned versions.
@@
For the Hugging Face models, we used the TNER library Ushio and Camacho-Collados (2021) to fine-tune each model, with entity decoding managed by CRF Lafferty et al. (2001). Hyperparameter tuning involved a grid search with learning rates of $10^{-4}$ and $10^{-5}$, both with and without weight decay set to 0.01. Training was carried out over 10 epochs, with an interim evaluation at 5 epochs, a batch size of 16, and gradient accumulation steps set to 2. Additionally, configurations included both CRF and non-CRF models, with a warmup step ratio of 0.1.
@@
Lastly, for GliNER, we fine-tuned the large model (“urchade/gliner_large”) on each dataset using customized hyperparameters. The main model was trained with a learning rate of $5\times 10^{-6}$ and weight decay set to 0.01. We used a linear learning rate scheduler with a warm-up ratio of 0.1. Training was carried out over 10 epochs with a batch size of 16, and focal loss parameters were set to alpha = 0.75 and gamma = 2. Step-based evaluations were performed every 300 steps.
@@ 10.3 Evaluation and statistical analysis
11 Results and discussion
Table 5 provides a comparison of the selected NER frameworks based on their macro-averaged F1 scores across ten datasets. The best and second-best performances are highlighted in bold and underlined, respectively. Our experiments reveal key performance trends influenced by dataset size, domain specificity, and the comparative effectiveness of pre-trained versus fine-tuned NER models. In the following, we delve into the details of these results and their implications for model performance across various datasets.
@@ 11 Results and discussion
Starting with large general-domain datasets such as CoNLL-2003, OntoNotes, Few-NERD, and MultiCoNER, Transformer-based models such as GliNER-L (fine-tuned), RoBERTa, and DeBERTa demonstrate outstanding performance. This can probably be attributed to their high parameter counts, which enable them to capture complex data relationships when trained on extensive datasets. A key insight from this analysis is the notable impact of GliNER-L’s comprehensive pre-training and fine-tuning processes. Compared to DeBERTa from Microsoft, GliNER-L exhibits improved performance, especially on domain-specific datasets. This improvement is likely due to GliNER-L’s structured approach, which involves sequential stages: pre-training as a language model, followed by NER training, and fine-tuning on domain-specific datasets. These stages allow GliNER-L to capture nuanced entity relationships and effectively adapt to NER tasks.
Shifting the focus to domain-specific datasets, particularly in the biomedical field (BioNLP2004, NCBI Disease, and BC5CDR), a different trend emerges. Traditional models such as CRF and LSTM-CRF remain competitive, often closely matching or even surpassing Transformer-based models. This may be due to the specialized terminology in biomedical texts, which is not well represented in the general datasets typically used for Transformer pre-training. However, fine-tuning on domain-specific data, as demonstrated with GliNER-L, significantly enhances performance by refining embeddings to better capture the contextual nuances of specialized terms. This trend is illustrated in Table 6, which compares the performance of CRF, LSTM-CRF, GliNER-L (fine-tuned) and Microsoft’s DeBERTa on key types of entity in biomedical datasets.
@@
In the BioNLP2004 dataset, which covers a broader range of entity types (“cell_line”, “cell_type”, “DNA”, “Protein”, and “RNA”), CRF consistently achieves high scores, especially in precision across entities like “cell_type” (94.54), “DNA” (89.32), “Protein” (93.96), and “RNA” (93.0). For “Protein” and “RNA” entities, GliNER-L shows competitive recall values (87.13 and 81.36, respectively), indicating its strength in capturing complex biomedical terms despite not leading in all F1 scores. LSTM-CRF and DeBERTa display relatively balanced scores across categories, though GliNER-L has a slight edge in recall for multi-word biomedical terms.
@@
Turning to smaller datasets such as FIN and WNUT2017, we observe that LSTM-CRF models tend to outperform Transformer-based models. This difference may stem from the tendency of overparameterized models, such as Transformers, to overfit when trained on limited data. For example, on the FIN dataset, LSTM-CRF achieves the highest F1 score (69.22), significantly outperforming Transformer models, which require careful hyperparameter tuning and well-structured validation sets to reduce overfitting. Notably, GliNER-L fine-tuned remains competitive even on these smaller datasets, consistently ranking among the top three models. This suggests that GliNER-L’s structured pre-training and fine-tuning processes enhance its robustness and adaptability across datasets of varying sizes.
@@
  • Precision/recall balance: Models located near the central line, such as DeBERTa variants and LSTM-CRF, exhibit a well-balanced trade-off between precision and recall. This balance makes them particularly suitable for tasks where an equilibrium between these metrics is crucial, ensuring both high precision and coverage in entity detection.
  • Precision-focused models: Models positioned above the reference line, such as CRF and CNNs (“en_core_web_sm” and “en_core_web_lg”), demonstrate higher precision than recall, making them ideal for applications where minimizing false positives is a priority.
  • Recall-focused model: GliNER-L pre-trained model favors recall, demonstrating considerably higher recall than precision, though it still falls significantly behind the other models in overall performance.
@@
Furthermore, Figure 9 presents a PCA plot (left) and hierarchical clustering (right), which illustrate performance similarities between frameworks. In the PCA plot, each point represents a model, with colors indicating clusters based on performance similarity. The first two dimensions capture most of the variance, showing that Transformer-based models, such as DeBERTa (base and large) and GliNER-L fine-tuned, cluster closely, suggesting consistent performance across datasets. In contrast, isolated points such as GPT-4o and GliNER-L pre-trained exhibit lower performance, particularly without fine-tuning. Hierarchical clustering (right) reinforces these insights. Traditional models, such as CRF and CNNs (both “en_core_web_sm” and “en_core_web_lg”), form a distinct cluster, highlighting their competitiveness but limited flexibility compared to Transformers. Fine-tuned Transformers, especially GliNER-L, stand out, underscoring the advantages of domain-specific training.
Statistical analysis supports these findings. A Friedman test applied to all entity types in our datasets reveals statistically significant differences between model medians (p-value = 0.000), confirming the significance at an alpha level of 5%. The results of Nemenyi’s post hoc test (Figure 10) show that certain groups of models have similar performance levels, as indicated by the connecting horizontal lines. For example, models such as GliNER-L fine-tuned and DeBERTa (large) achieve the best rankings, indicating superior performance, and are not significantly different from each other. This suggests that these models perform similarly on the evaluated datasets. In particular, among the Transformer-based models, the DeBERTa variants appear to be the most effective for NER.
Additionally, another cluster of models, including GliNER-L pre-trained, CNNs (“en_core_web_sm” and “en_core_web_lg”), and GPT-4o, also exhibit statistically similar performance, although with higher mean ranks, indicating relatively lower performance compared to the top performing models. Toward the right of the figure, models such as CRF, DistilBERT, and BERT exhibit the highest ranks, indicating inferior performance in comparison to the other models.
@@
Our experiments have shown that while LLMs perform exceptionally well in many NLP tasks, their general nature often leads to a lack of specialized effectiveness for NER, which can impede precise entity recognition, particularly in noisy or dynamic contexts. Evaluations highlight a performance gap for NER tasks Han et al. (2023), likely due to limited task-specific learning and explicit understanding, which may lead to a “lack of specialty” in NER. Moreover, the large, sometimes undisclosed, number of parameters in LLMs poses challenges for fine-tuning to meet the demands of NER. Consequently, directly addressing NER tasks using only LLMs remains difficult. However, combining LLMs with simpler and fine-tuned NER models offers a promising solution. Approaches such as LinkNER Zhang et al. (2024b) leverage a hybrid model in which the NER component handles common entities, while the LLM, guided by uncertainty estimation, addresses complex or ambiguous cases, enhancing robustness and flexibility in open domain contexts. Similarly, the Super In-Context Learning (SuperICL) approach Xu et al. (2023) integrates smaller specialized plug-in models to provide task-specific predictions and confidence scores, refining LLM outputs for more accurate final predictions. Together, LinkNER and SuperICL demonstrate that combining the broad contextual knowledge of LLMs with the focused precision of smaller models offers a promising path to advance NER, particularly in challenging open-domain settings with unseen entities and noisy data.
@@ 12 Conclusion and perspectives
This article presents a comprehensive survey on recent advances in NER within a classification framework. We focus on recent methods, including LLMs, graph-based approaches, reinforcement learning techniques, and strategies to train models on small datasets. To assess these approaches, we evaluated popular frameworks across datasets with varying characteristics.
@@
Looking ahead, the potential of LLMs to enhance NER should not be underestimated. Despite current limitations in handling specialized NER tasks, the rapid advancement of language models presents an exciting opportunity to integrate these technologies into more refined NER systems. Promising directions to advance NER include hybrid approaches that combine LLMs with specialized, fine-tuned NER models (such as LinkNER or SuperICL). By blending the broad contextual understanding of LLMs with the precision of specialized NER models, these strategies hold promise to tackle NER in dynamic, open-domain settings that involve unseen entities and noisy data. Future research should further investigate these hybrid approaches, along with fine-tuning techniques and preprocessing methods designed to enhance LLM adaptability for NER. Such strategies could improve the applicability and accuracy of NER systems across diverse datasets and challenging contexts.